in a exam there are 90 mcq’s.for each correct answer 4 marks are awarded.for each incorrect answer 2 marks are deducted.and in case of unattempted question 0.5 are deducted
a student left 10 question and scores a total of 195 marks.tell how many questions does he correct?

Grade:8

1 Answers

Arun
6 years ago
You can check by hit and trial method
Otherwise propwr method is given below
That he attempted 80 questions
Suppose correct questions = x
Then incorrect = 80 -x
Now
4x - 2(80-x) -0.5*10 = 195
6x -160 -5 = 195
6x = 360
X = 60
Hence correct questions = 60